### v1

# DualReal: Adaptive Joint Training for Lossless Identity-Motion Fusion in Video Customization

Abstract
Customized text-to-video generation with pre-trained large-scale models has recently garnered significant attention through focusing on identity and motion consistency. Existing works typically follow the isolated customized paradigm, where the subject identity or motion dynamics are customized exclusively. However, this paradigm completely ignores the intrinsic mutual constraints and synergistic interdependencies between identity and motion, resulting in identity-motion conflicts throughout the generation process that systematically degrades. To address this, we introduce DualReal, a novel framework that, employs adaptive joint training to collaboratively construct interdependencies between dimensions. Specifically, DualReal is composed of two units: (1) Dual-aware Adaptation dynamically selects a training phase (i.e., identity or motion), learns the current information guided by the frozen dimension prior, and employs a regularization strategy to avoid knowledge leakage; (2) StageBlender Controller leverages the denoising stages and Diffusion Transformer depths to guide different dimensions with adaptive granularity, avoiding conflicts at various stages and ultimately achieving lossless fusion of identity and motion patterns. We constructed a more comprehensive benchmark than existing methods. The experimental results show that DualReal improves CLIP-I and DINO-I metrics by 21.7% and 31.8% on average, and achieves top performance on nearly all motion quality metrics. Project page: https://wenc-k.github.io/dualreal/
Figure 1: Generated customization results of our proposed novel paradigm DualReal. Given subject images and motion videos, DualReal generates high-quality customized identity and motion simultaneously, without compromising the consistency of either dimension.
1 Introduction
Video constitutes a spatiotemporal embodiment of the real world, where the spatial subject identity and the temporal motion dynamics form mutually constrained yet synergistic dimensions for physical modeling. This mutuality manifests through their inherent interdependence, i.e., maintaining stable subject identities across frames restricts motion possibilities, while enforcing certain motion trajectories conversely necessitates corresponding topological transformations of identity representation (e.g., 180° viewpoint rotation leads to identity transformation from frontal to dorsal profiles).
Video customized generation [21, 46, 54, 49, 48], which aims to mimic the user-specified concepts (i.e., subject identities, dynamic motions, or both) beyond linguistic expressibility, significantly enhances the controllability of video synthesis systems. This task greatly expands the applicability scope of pre-trained text-conditioned video models to cinematic production, personalized avatars, etc., attracting growing interest from academic and industrial communities. The primary challenge of customized video generation lies in two interdependent dimensional objectives, i.e., (1) identity consistency, i.e., the target subject should closely match the given reference in all frames, while minimizing temporal motion artifacts, and (2) motion consistency, i.e., the subject motion should closely match the given reference across frames, while minimizing spatial identity artifacts.
As a rapidly emerging research frontier, existing video customized methodologies currently focus on either identity or motion independently. VideoBooth [21] achieves subject identity-driven generation by injection of reference image embedding, while AnimateDiff [11] achieves the animation of static outputs into videos by appending trainable temporal modules to personalized text-to-image models. Recently, DreamVideo [48] employs independent training for each dimension (i.e., identity or motion) and directly blends their parameters during inference to achieve both the identity and motion customization simultaneously, demonstrating promising results to combine specific subject identity and motion patterns. Essentially, current works typically follow the isolated customized paradigm, where the subject identity or motion dynamics are customized exclusively.
However, the existing isolated customized paradigm completely ignores the intrinsic mutual constraints and synergistic interdependencies between identity and motion, resulting in identity-motion conflicts throughout the generation process that systematically degrades either motion coherence, subject fidelity, or both dimensions simultaneously. The reason is that the step-by-step diffusion video synthesis process itself, by nature, with different denoising steps dynamically reweights their spatiotemporal focus, i.e., progressively refines identity details across frames, enabling completely modeling through increasing denoising steps. The existing isolated customized paradigm violates this natural progression through indiscriminate dimensional over-specialize across all time steps, since they enforce uniform step sampling during motion/identity customized training, thereby inducing conflicting optimization trajectories between motion and identity accuracy. Consequently, existing methodologies inevitably cause mutual performance deterioration, that is, motion customization undermines the pre-trained video model’s inherent identity priors, or vice verse. As shown in Fig. 2, (1) adding the motion prior irreversibly reduced identity fidelity, indicating that isolated training fails to resolve dimensional conflicts during inference and leads to performance degradation; and (2) as motion training steps increased, optimal fidelity for different identities occurred unpredictably, suggesting that no universal number of training steps minimizes degradation. In summary, the existing paradigm does not meet the consistency and flexibility requirements of video customization tasks.
Figure 2: Visual analysis of isolated training paradigm. We select different identities with the same motion pattern, fix the number of identity training steps, and gradually increase the number of motion training steps to achieve two-dimension customization. The red box marking the relative optimal position of the same identity’s fidelity. Experiments show that (1) adding motion prior significantly damages identity consistency; (2) We cannot find a universal step to minimize identity degradation from the positions of the red boxes for different identities.
In this paper, we introduce DualReal, a novel framework that, for the first time, employs adaptive joint training to collaboratively construct interdependencies between identity and motion, which meets the consistency requirements of customized video generation in both identity and motion, as shown in Fig. 1. Technically, DualReal is composed of two complementary units: (1) Dual-aware Adaptation dynamically selects the appropriate training phase(i.e., either identity or motion) to learn the current information, guided by the other dimension prior. It also employs a regularization strategy to prevent dimensional knowledge leakage by blocking the parameter updates of the non-training dimension using gradient masking, thereby achieving effective joint training; (2) StageBlender Controller operates through coordinated utilization of denoising-stage progression and Diffusion Transformer (DiT) layer-depth variations during training. By adaptively allocating hierarchical focus (i.e., fine-grained adjustments for identity patterns and motion dynamics), it resolves dimensional competition across processing stages. This granularity-aware guidance ultimately achieves dimensional lossless fusion.
Our contributions can be summarized as follows:
Concepts. For the first time, we (1) point out that the isolated paradigm causes mutual performance deterioration (identity fidelity and motion coherence) because it ignores the intrinsic constraints and synergistic; (2) present DualReal, a novel paradigm that employs adaptive joint training to collaboratively construct interdependencies.
Technology. The proposed DualReal framework consists of two components: (1) Dual-aware Adaptation, which alternates between identity and motion training phases, leveraging dimension-specific guidance and regularization to prevent information leakage and enable joint training; (2) StageBlender Controller, which adaptively coordinates denoising stages and DiT depths to guide modes at different granularities, resolving conflicts and enabling seamless fusion of identity and motion patterns.
Performance We constructed a more comprehensive evaluation benchmark than existing methods. The experimental results show that DualReal improves CLIP-I and DINO-I metrics by 21.7% and 31.8% on average, and achieves top performance on nearly all motion quality metrics, demonstrating the efficiency of our adaptive joint training framework.
2 Related Work
2.1 Text-to-video Diffusion Models
Text-to-video generation aims to generate realistic videos based on prompts and has recently received growing attention [8, 1, 53, 30, 15, 44, 38, 11, 50, 52, 55, 45, 42, 28]. Current text-to-video generation architectures primarily fall into two categories[33, 29]. UNet-based video diffusion frameworks utilize hierarchical enc-dec with spatiotemporal learning [38, 11, 16, 3, 50, 44, 15], e.g., Video diffusion models [15] pioneered diffusion model applications in video generation through pixel-space video distribution modeling. Make-A-Video [38] and AnimateDiff [11] augment pretrained text-to-image models with motion modules. Recent advances in scalability drive the shift toward transformer-based architectures with joint spatiotemporal modeling [28, 52, 20, 26, 22], achieving revolutionary progress in video generation. Sora [28] introduces the diffusion-transformer framework, achieving cinematic-quality extended video synthesis with temporal stability. CogVideoX [52] introduces an expert transformer with adaptive LayerNorm for enhanced text-video feature fusion. While diffusion transformers exhibit strong generative capacities, their architectural constraints in spatiotemporal decoupling inherently limit dynamic concept embedding.
2.2 Generation Model Customization
In contrast to domain-agnostic generation frameworks, customized visual synthesis exhibits superior adaptability in addressing personalized visual requirements via parametric adaptation mechanisms[6, 18, 7, 12, 34, 35, 37, 39, 47, 9]. Textual inversion [9] aligns visual-textual semantics through text embedding optimization. Dreambooth [34] through full model fine-tuning of diffusion architectures to inject subject-specific priors. Building upon these foundational approaches, contemporary video customization research explores analogous methodologies[27, 5, 25, 13, 48, 49, 32, 2, 10]. MotionBooth [49] proposes comprehensive video diffusion model fine-tuning coupled with attention map manipulation for motion control during inference. DreamVideo [48] develops decoupled adapter training with joint inference mechanisms, coordinating subject customization and motion preservation during generation.
3 Methodology
Figure 3: Overall framework of DualReal. At each training step, we first dynamically select the training phase Z(i.e., identity or motion) to determine the data processing path. The specific data undergoes noise injection and combines with the text embeddings. StageBlender Controller governs two-dimension adapters’ contributions in Dual-Aware Block (DA-Block) through time-aware conditioning of current denoising step and fused feature representations. In DA-Block, the training-stage(i.e. Z) adapter learns the current information guided by the frozen dimension prior, and employs a regularization strategy to avoid dimensional knowledge leakage, achieving joint training. Both branches engage in residual connections with DiT outputs.
Given a series of identity-specific images and motion sequences, DualReal through an innovative joint training framework, synthesizes coherent motion while preserving full-frame identity fidelity, as detailed in Fig. 3. During training, DualReal dynamically switches optimization focus between identity and motion, adjusting the module parameters through specific guidance. To address the key challenges: (1) enabling joint identity-motion interaction modeling in unified parameter spaces. (2) mitigating attribute leakage risks in alternating training modes. We propose the Dual-aware Adaptation architecture with a complementary regularization strategy in Sec. 3.2, forming an integrated framework for parameter-shared adaptation. Moreover, we propose StageBlender Controller, which leverages the denoising stages and DiT depths to guide different modes with adaptive granularity, avoiding dimensional conflicts at various stages and ultimately achieving lossless fusion of identity and motion in Fig. 5.
3.1 Preliminary
DiT-based Models. Most DiT-based models [52, 28, 22, 26] process concatenated conditioning prompt and spatiotemporal visual tokens through transformer layers, establishing multi-modal coupling between text-guided semantic contexts and visual representations in latent space. Despite achieving a remarkable capabilities in generic generation scenarios, this architectural fundamentally constrains conventional personalization frameworks that demand decoupled control along spatial-temporal axes [46, 49].
3.2 Dual-aware Adaptation
To achieve joint training of identity and motion and resolve dimensional conflicts, we innovatively proposed Dual-aware Adaptation, which leverages the prior from one dimension to guide the training of the other while preventing information leakage through a regularization strategy, as shown in the lower half of Fig. 3.
Joint Identity-Motion Optimization. Different from some approaches that fine-tune the whole diffusion model  [34], DualReal first dynamically select the training dimension (i.e., motion-focused or identity-focused) with the predefined hyperparameter ratio before each denoising iteration, the corresponding data is then sent to the DiT. The input of $i$-th block is the joint feature $f_{\text{in}}^{i}=[f_{\text{text}}^{i},f_{\text{viusal}}^{i}]\in\mathbb{R}^{B\times(n_{t}+n_{v})\times c}$, where $n_{t}$ , $n_{v}$ represent the number of text and visual tokens respectively, adapters employ the bottleneck architecture with skip connections  [48], which can be expressed as:
$f_{\text{id}}^{i}=\sigma\left(f_{\text{in}}^{i}*\mathbf{W}_{\text{down}}*\mathbf{W}_{\text{up}}\right),$ (1)
$f_{\text{mo}}^{i}=\sigma\left((f_{\text{in}}^{i}*\mathbf{W}_{\text{cond}})*\mathbf{W}_{\text{down}}^{{}^{\prime}}*\mathbf{W}_{\text{up}}^{{}^{\prime}}\right),$ (2)
where the activation function $\sigma$ corresponds to GELU  [14], $\mathbf{W}$ and $\mathbf{W^{{}^{\prime}}}$denote the identity-preserving and motion-specific linear projection weights respectively, both operating on hidden dimension $d$. The weight $\mathbf{W}_{\text{{cond}}}\in\mathbb{R}^{e\times c}$ of conditional linear maps reference image embedding to the latent space [48].
Through StageBlender Controller constraint (Fig. 5) , the motion adapter outputs are scaled by weight coefficient $\omega_{i}$ , with identity outputs weighted by the complementary coefficient ($1-\omega_{i}$). The modulated features are aggregated into the output of DiT blocks through residual connections, above process can be formulated as:
$\hat{f}_{\text{out}}^{i}=\omega_{i}*f_{\text{mo}}^{i}+(1-\omega_{i})*f_{\text{id}}^{i}+f_{\text{dit}}^{i}\ ,$ (3)
where $f_{\text{dit}}^{i}$ denoting the output of the $i$-th DiT layer and $\hat{f}_{\text{out}}^{i}$ indicating the aggregated output of final block. This parametric constraint intrinsically balances feature contributions across blocks and denoising stages through complementary scaling, while structurally enforcing dedicated attention to either identity preservation or motion dynamics during adaptation.
Regularization Strategy. A critical challenge in joint dimensional training arises from the significant distribution shift across different training dimension, where unconstrained optimization usually causes destructive interference between cross-dimension knowledge observed in previous work [51, 49]. For example, fine-tuning the motion adapter with static images during the temporal training phase irreversibly degrades its dynamic generation capability, with analogous effects occurring during identity adaptation. To resolve this, we employ the Regularization with the gradient mask M, to activate only the corresponding adapter parameters based on the selected stage Z, optimizing motion coherence(i.e., through motion adapter), or preserving identity consistency(i.e., through identity adapter), which can be formulated as:
$\theta^{(t+1)}=\theta^{(t)}-\bf M\odot\nabla_{\theta}\mathcal{L}\ ,$ (4)
${\bf M}={\bf Z}\cdot M_{m}+(1-{\bf Z})\cdot M_{i}\ .$ (5)
The mask conditions are defined as:
$\begin{cases}M_{m}[l]=1\iff M_{m}\cdot\theta[l]=\theta_{m}\ ,\\ M_{i}[k]=1\iff M_{i}\cdot\theta[k]=\theta_{i}\ ,\end{cases}$ (6)
where $\mathcal{L}$ denotes the video diffusion reconstruction loss, Z is the binary selector, and $\theta_{m}/\theta_{i}$ represent the motion/identity adapter parameters respectively.
Simultaneously, we keep frozen adapter in a waiting state to inform modal expertise for active adapter during forward propagation, enabling two dimensional features referencing within current data streams. The features from frozen adapter act as intrinsic regularization to constrain dimension overfitting, thereby facilitating mutual reference learning without interference.
3.3 StageBlender Controller
Figure 4: Illustration of proposed StageBlender Controller, which employs Adaptive LayerNorm mechanism modulates text-visual feature based on timestep-conditional embeddings, then maps the feature to multiple groups after residual gated connections. These scaled weight are subsequently routed to their respective DA-Blocks for processing.
Figure 5: Qualitative comparison with existing methods. Compared with other methods, DualReal achieves high identity consistency with coherent motion, demonstrating the advantage of joint training in balancing pattern conflicts.
Furthermore, in order to resolves dimensional competition across processing stage, we propose the StageBlender Controller that governs dimensional contributions through time-aware conditioning of block-level scaling coefficients, which empowers the DA-Block to adaptively allocate specific dimension shares (i.e., achieve granularity decoupling) through the mechanism detailed in Fig. 4. Specifically, this module dynamically generates multiple sets of scaling weights according to denoising timestep embedding and the fused text-visual features. For the input feature $f_{\text{in}}=[f_{\text{text}}^{1},f_{\text{viusal}}^{1}]\in\mathbb{R}^{B\times(n_{t}+n_{v})\times c}$, the processing flow first extracts salient features through pooling, then adaptively modulates them via DiT Adaptive LayerNorm  [52] with injected timestep embeddings $\bf t$. This operation can be formulated as:
$f^{{}^{\prime}}=\text{{Pooling}}(f_{\text{in}},\ \text{dim=1})*\mathbf{W}\ ,$ (7)
$f^{{}^{\prime\prime}}=\text{MLP}\left(\text{LaynerNorm}(f^{\prime})\right)*\mathbf{\alpha}+\mathbf{\beta}\ ,$ (8)
where the $\mathbf{W}\in\mathbb{R}^{c\times\mathbf{t}_{\text{dim}}}$ denotes the weight matrix with $\mathbf{t}_{\text{dim}}$ as the channel dimension of timestep embedding, and $\alpha$, $\beta$ are defined as:
$\mathbf{h}=\text{MLP}\left(\text{SiLU}(\mathbf{t})\right)\ ,$ (9)
$\mathbf{\alpha},\ \mathbf{\beta},\ \mathbf{\gamma}=\mathbf{h}_{:d},\mathbf{h}_{d:2d},\mathbf{h}_{2d:3d}\ .$ (10)
The computed weights coefficients are then integrated to enable gated fusion between timestep and visual text tokens, as formulated below:
$f_{g}=f^{{}^{\prime\prime}}+\mathbf{\gamma}*f^{{}^{\prime}}.$ (11)
Through empirical analysis of DiT-based denoising architectures, we observe that deeper blocks inherently specialize in processing concrete, fine-grained features. To enhance hierarchical decoupling, we implement a downward-propagating MLP that transforms integrated features into weight groups, as formalized below:
$\underbrace{\mathbf{\omega}^{(1)},...,\mathbf{\omega}^{(n)}}_{\text{Weights groups}}=\ \text{softmax}\left(\mathbf{\Gamma}\cdot(\text{MLP}(f_{g}))\right)\ ,$ (12)
where $\mathbf{L}$ corresponds to the depth of the transformer block, $\mathbf{n}$ specifies the cardinality of the disentangled weight groups and $\mathbf{\Gamma}$ is the projection operator: $\mathbb{R}^{\mathbf{L}}\to\mathbb{R}^{\mathbf{n}}$. Each group sequentially governs designated DA-Block layers through parameter assignment.
4 Experiment
Method Text Consistency Identity Similarity Motion Quality
CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO-I $\uparrow$ T.Fli $\uparrow$ T.Cons $\uparrow$ MS $\uparrow$ Dynamic Degree
MotionBooth [49] 0.317 0.566 0.459 0.962 0.972 0.973 10.95 ( -1.07 )
LoRA [17] 0.323 0.425 0.286 0.956 0.976 0.973 25.34 ( +13.32 )
CogVideoX-5B [52] 0.336 0.521 0.424 0.947 0.973 0.965 26.51 ( +14.49 )
DreamVideo [48] 0.278 0.458 0.334 0.949 0.963 0.968 8.841 ( -3.18 )
DualReal (Ours) 0.323 0.629 0.551 0.965 0.983 0.978 14.96 ( +2.94 )
Table 1: Quantitative comparison of personalization video generation for customized subject and motion. We highlight the best and second-best values for each metric. “T.Cons” and “T.Flickering” denotes Temporal Consistency and Temporal Flickering, respectively. Compared with other methods, DualReal achieved average improvements of 21.7% on CLIP-I and 31.8% on DINO-I, recorded the best results on three motion quality metrics (T.Cons, Motion Smoothness(MS), and Temporal Flickering), and ranked second on CLIP-T. The motion datasets achieve an average Dynamic Degree of 12.02 and parenthetical values quantify the current method’s deviation from this benchmark to determine the intensity consistency of movement.
4.1 Setup
Datasets. The evaluation datasets are divided into two components: identity images and motion videos. For identity customization, 50 subjects are strictly selected from previous works  [23, 24] and Internet collections (including pet, plush, etc.), with each subject containing 3–10 images. For motion customization, 21 motion sequences with challenging dynamic patterns are collected from public datasets [41, 40]. Additionally, each case is provided with 50 various prompts containing different editability (i.e., decoration or environment) to sequentially evaluate the method’s editability and scene versatility.
Baselines. Among existing methods, only DreamVideo [48] achieves customization of both subject and motion. For fair comparison, we implement two approaches from the same backbone: (1) CogVideoX-5B [52]: Sequential full-parameter fine-tuning with identity then motion data as in DreamBooth [34] paradigm. (2) LoRA fine-tuning [17]: Separate training of two LoRA modules for identity and motion, then fuse their parameters during inference. Additionally, the identity module of MotionBooth [49] introduces irrelevant random videos during training to preserve the model’s motion capability, so we compare our approach with this method as well. In summary, we evaluate our results against DreamVideo, CogVideoX-5B, LoRA fine-tuning and MotionBooth to provide a more comprehensive performance analysis.
Evaluation metrics. We use seven metrics across three dimensions. (1) Text-Video Consistency is measured by CLIP-T scores, computed as the CLIP [31] cosine similarity between text prompts and all generated frames. (2) Identity Fidelity is quantified using DINO-I and CLIP-I scores, which assess feature similarity between generated frames and reference identity images via DINO ViTS/16 [4] and enhanced CLIP [36] embeddings, respectively. (3) Temporal Motion Quality is evaluated with four metrics: T-Cons [8] for temporal consistency, Motion Smoothness (MS) for global fluidity, Temporal Flickering (TF) for high-frequency inconsistencies measured by mean absolute differences between adjacent frames, and Dynamic Degree (DD) leveraging RAFT optical flow estimation [43] to quantify motion intensity (We quantify the method’s deviation from the benchmark to determine the intensity consistency of movement). Notably, MS, TF, and DD are adopted from the comprehensive video benchmark VBench [19].
4.2 Main Results
Qualitative results. Qualitative experiments in Fig. 5 show that while MotionBooth maintains identity fidelity, it fails to model motion patterns effectively. DreamVideo suffers from pattern conflicts during inference, resulting in inconsistent identity. Similarly, CogVideoX-5B and LoRA struggle to preserve identity due to their decoupled training methods. In contrast, DualReal achieves high identity consistency with coherent motion, demonstrating the advantage of joint training in balancing pattern conflicts.
Quantitative results. As shown in Tab. 1, DualReal achieved average improvements of 21.7% on CLIP-I and 31.8% on DINO-I, recorded the best results on three motion quality metrics (T.Cons, Motion Smoothness, and Temporal Flickering), and ranked second on CLIP-T. Although our DD metric for quantifying motion intensity is not high, we evaluate all motion data and found an average DD of 12.02. Our metric deviates slightly from it, proving there is no collapse in motion amplitude. Overall, our method significantly enhances motion coherence and identity fidelity while preserving text consistency, further validating our adaptive joint training approach.
4.3 Ablation Studies
We evaluate our method by conducting ablation studies on a smaller evaluation subset, with the observed trends aligning with those of the main evaluation set. Additional ablation results are provided in the supplementary.
Settings CLIP-T CLIP-I DINO-I DD
w/o Dual-aware Adaptation 0.334 0.616 0.647 3.51(-5.53)
w/o StageBlender Controller 0.346 0.619 0.652 5.70(-3.31)
w/o Weight Groups 0.335 0.662 0.766 5.83(-3.12)
ours 0.333 0.674 0.771 6.34(-2.70)
Table 2: Quantitative ablation studies on each component. We implement Dual-aware Adaptation removal by separately training the two modalities and directly blends their parameters during inference, following the approach of DreamVideo. The motion datasets achieve an average Dynamic Degree of 9.04.
Figure 6: Qualitative ablations studies on each component. (1) Omitting Dual-aware Adaptation introduces artifacts on the subject’s hands and chin, significantly reducing clarity. Using fixed weights for the dimensional adapters without the StageBlender Controller causes the hands to become overly adapted to the motion pattern, and removing weight grouping reduces identity fidelity and background detail.
Group Cardinality CLIP-T CLIP-I DINO-I DD
n=1 0.335 0.662 0.766 5.83 (-3.21)
n=2 0.343 0.632 0.660 5.49 (-3.55)
n=42 0.336 0.631 0.706 6.24 (-2.80)
ours(n=7) 0.333 0.674 0.771 6.34 (-2.70)
Table 3: Quantitative ablation studies of group cardinality. The results suggest that very small groups may lack sufficient context and overly large groups may dilute crucial details, making balanced group cardinality essential for optimal performance.
Quantitative experiment on each component. The ablation study in Tab. 2 shows that removing Dual-aware Adaptation or the StageBlender Controller slightly increases the text consistency metric CLIP-T but significantly decreases identity similarity and motion intensity, highlighting the need for joint dimension training and granular control. Additionally, metric changes diminish slightly when weight groups are removed.
Qualitative experiment on each component. Qualitative results in Fig. 6 show that omitting Dual-aware Adaptation introduces artifacts on the subject’s hands and chin, significantly reducing clarity. Using fixed weights for the dimensional adapters without the StageBlender Controller (i.e., direct fusion at inference) causes the hands to become overly adapted to the motion pattern. Removing weight grouping (i.e., uniformly modulating all blocks) reduces identity fidelity and background detail compared to our final configuration. These findings confirm that each component is essential for high-quality customization generation.
Effectiveness of group cardinality. As shown in Tab. 3, CLIP-I and DINO-I performance declines when group cardinality is either very small or very large, while a balanced group size (n=7) yields the best results. This suggests that very small groups may lack sufficient context and overly large groups may dilute crucial details, making balanced group cardinality essential for optimal performance.
4.4 Visual analysis of StageBlender Controller
As Fig. 7 shows, shallow blocks (Groups 1–6, blue) progressively increase identity weights during denoising, emphasizing early identity preservation. In contrast, the deepest block (Group 7, red) steadily raises motion weights to enhance motion modeling. Overall (orange dashed line), as denoising advances, the model increasingly prioritizes identity preservation over motion generation, highlighting the distinct roles of different network depths.
These observations further confirm that: (1) as denoising progresses, the diffusion model’s emphasis shifts monotonically between identity and motion, with a growing focus on identity; and (2) DiT networks of different depths divide the tasks of modeling identity and motion differently at each denoising step, with the deepest network focusing on motion patterns and showing increased enhancement as denoising advances.
Figure 7: Controller Visual Analysis. We show the Identity Weights trends across denoising steps for different block depths. (1) As denoising progresses, the diffusion model’s emphasis shifts monotonically between identity and motion, with a growing focus on identity(orange dashed line); (2) The deepest block group exhibits an inverse pattern, i.e., with the denoising process increasingly prioritizing motion coherence modeling.
5 Conclusion
In this paper, we propose DualReal, a novel approach for customized video generation given a subject and motion. DualReal adaptively trains identity and motion jointly, resolving dimensional conflicts and enabling universal sample customization. Our framework leverages the prior from one dimension to guide the training of the other, while preventing information leakage through a regularization strategy. Simultaneously, we use a controller to guide the lossless fusion of modes based on various denoising stages and DiT depths. Evaluated on a more comprehensive evaluation benchmark, our method improves CLIP-I and DINO-I metrics by 21.7% and 31.8% on average, and achieves top performance on nearly all motion quality metrics, demonstrating the efficiency of our adaptive joint training framework.